1. Find the PV of an annuity consisting of 60 quarterly payments of Rs. 200, the first being made at the end of 4 years 9 months. Assume that the rate of interest is 12.3% p.a. compounded monthly.
1
Expert's answer
2021-05-02T07:33:08-0400

We have to calculate the PV of annuity that starts at time t = 4 years 9 months and then discount this back to t=0 to get the value as of today.

"PV of annuity = PMT \times\frac{1 − (1 + i)^{−n}}{i}"

PV=present value

PMT=amount in each annuity payment

I=interest

n=number of payment


Step 1

Effective Annual Rats of the rate of interest:

12.3% compounded monthly

"= (1+\frac{r}{m})^{m}-1"


"= (1+\frac{0.123}{12})^{12}-1"


"=0.130176595 = 13.02" %

Value of annuity at 4 years 9 months from today:

"PV = 200\times[\frac{1−(1+\frac{0.1302}{4})^{-60}}{\frac{0.1302}{4}}]"

"PV=5245.91901"


Step 2

We need to discount this value of $5,245.92 to today's value,

4 years and 9 months is "4\times4 +3 = 19" quarters

"PV0=\frac{ PV1}{(1+r)^{19}}"


"=\frac{ 5,245.92 }{(1+\frac{0..1302}{4})^{19}}"


"=2,854.68"

The present value of the annuity today is $2,854.68